OpenStudy (anonymous):

Prove that: √e >= 3/2 . Any options for this inequality?

OpenStudy (mimi_x3):

Is it:\[\large \sqrt{e} \geq\frac{3}{2} \]??

OpenStudy (anonymous):

yes. that's it.

OpenStudy (mimi_x3):

And is \(e\) \(2.7...\) or is it just a variable?

OpenStudy (anonymous):

e is the constant also.

OpenStudy (mimi_x3):

so its a variable like \(x\)?

OpenStudy (anonymous):

no, is the Euler's number. 2,7 as you previously said.

OpenStudy (mimi_x3):

sorry i dont know..i should have read the question properly before answering :/ https://www.wolframalpha.com/input/?i=%E2%88%9Ae+%3E%3D+3%2F2